Required information

[The following information applies to the questions displayed below.]

Gabi Gram started The Gram Co., a new business that began operations on May 1. The Gram Co. completed the following transactions during its first month of operations.

May 1 G. Gram invested $43,000 cash in the company in exchange for its common stock.
1 The company rented a furnished office and paid $2,300 cash for May’s rent.
3 The company purchased $1,880 of office equipment on credit.
5 The company paid $730 cash for this month’s cleaning services.
8 The company provided consulting services for a client and immediately collected $5,100 cash.
12 The company provided $2,700 of consulting services for a client on credit.
15 The company paid $780 cash for an assistant’s salary for the first half of this month.
20 The company received $2,700 cash payment for the services provided on May 12.
22 The company provided $3,300 of consulting services on credit.
25 The company received $3,300 cash payment for the services provided on May 22.
26 The company paid $1,880 cash for the office equipment purchased on May 3.
27 The company purchased $85 of office equipment on credit.
28 The company paid $780 cash for an assistant’s salary for the second half of this month.
30 The company paid $350 cash for this month’s telephone bill.
30 The company paid $260 cash for this month’s utilities.
31 The company paid $1,800 cash in dividends to the owner (sole shareholder).

2-a. Prepare income statement for May.
2-b. Prepare statement of retained earnings for May.
2-c. Prepare balance sheet for May 31.
3. Prepare statement of cash flows for May.
Answer #1

Part 2 A

THE GRAM CO.

INCOME STATEMENT

For Month Ended May 31

Revenues

Consulting services revenue

11100

Expenses

Rent expense

2300

Salaries expense

1560

Cleaning expense

730

Telephone expense

350

Utilities expense

260

Total expense

-5200

Net income

5900

Consulting services revenue = 5100+2700+3300 = 11100

Salaries expense = 780*2 = 1560

Part 2 B

THE GRAM CO.

STATEMENT OF RETAINED EARNINGS

For Month Ended May 31

Retained earnings - May 1

0

net income

5900

5900

Dividends

-1800

Retained earnings - May 31

$4100

Part 2 C

THE GRAM CO.

BALANCE SHEET

May 31

Asset

Liabilities

Cash

45220

Accounts payable

85

Office equipment

1965

Equity

Common stock

43000

Retained earnings

4100

47185

47185

Office equipment = 85+1880 = 1965

Cash = 11100-2300-730-350-260-1560-1880+43000-1800 = 42780

Part 3

THE GRAM CO.

Statement of Cash Flows

For Month Ended May 31

Cash flows from operating activities

Cash received from customers

11100

Cash paid for rent

-2300

Cash paid for cleaning

-730

Cash paid for telephone

-350

Cash paid for utilities

-260

Cash paid to employees

-1560

Net cash provided by operating activities

5900

Cash flows from investing activities

Purchase of equipment

-1880

Net cash used by investing activities

-1880

Cash flows from financing activities

Investment by owner

43000

Withdrawal by owner

-1800

Net cash provided by financing activities

41200

Net increase in cash

45220

Cash balance, May 1

0

Cash balance, May 31

45220